EIGENSTATE RESOLVED INFRARED AND MILLIMETER WAVE - INFRARED DOUBLE RESONANCE STUDY OF METHYLAMINE IN THE N-H STRETCH OVERTONE REGION

\underline{H. K. SRIVASTAVA}, A. CALLEGARI, K. K. LEHMANN, G. SCOLES, Princeton University, Department of Chemistry, Princeton, NJ, 08544; U. MERKER^{a}, University of Bonn, Institut für Angewandte Physik, Wegelerstr. 8, 53115 Bonn, Germany.

The infrared spectrum of methylamine, CH_{3}NH_{2}, was recorded in the region of the ν_{1} + ν_{10} N-H stretching overtone (6600 cm^{−1}), using a molecular beam laser spectrometer with optothermal detection. This molecule is an interesting benchmark for the study of intramolecular vibrational energy redistribution (IVR) because of its two large amplitude internal motions: CH_{3} rotation and NH_{2} wagging that are coupled together. However this also makes assigning the spectrum a challenge due to its extreme congestion. The single photon spectrum yields a density of approximately 1000 lines/cm^{−1}. Millimeter wave-infrared double resonance has been used to isolate the contributions from a single ground state. Preliminary inspection of the double resonance spectrum suggests that the IVR rates are faster than those observed for molecules of comparable structure and size.^{b} Analysis of the data is in progress, and further results will be presented at the conference.

aSupported by the German Academic Exchange Service (DAAD)

bTalk RK11, 1995.